OpenStudy (anonymous):

Write a direct variation equation that relates x to y. Then solve. Show both the equation and the solution. If y = 8 when x = 4, find y when x = 5.

OpenStudy (ddcamp):

Direct variation means that there is some number k such that: y=k*x . Can you find k given y=8 when x=4?

OpenStudy (anonymous):

i dont think so im really new to this could u expain it to me???

OpenStudy (ddcamp):

So, we're given an initial condition: "y=8 when x=4," and we're told that we have a direct variation equation. This means that, given x, y=k*x. We plug the given values for x and y into the equation, giving us: 8=k*4 Can you solve for k?

OpenStudy (anonymous):

wait i think ive got it y=kx 8=k * 4 k=8/4=2 y=2x y=2 * 5 y=10

OpenStudy (anonymous):

right?

OpenStudy (ddcamp):

Yes

OpenStudy (anonymous):

awesome! thanks so much!